OpenStudy (anonymous):

2x^-1/4•2y^3/4 ------------------- x^2y^1/2

OpenStudy (teddyiswatshecallsme):

Instructions?

OpenStudy (anonymous):

I need to simplify

OpenStudy (teddyiswatshecallsme):

What level math is this?

OpenStudy (acxbox22):

can u draw the question out?

OpenStudy (anonymous):

Sure hold on a sec

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

lets write it using math

OpenStudy (anonymous):

Computer is slow lol

OpenStudy (misty1212):

\[\huge \frac{2x^{-\frac{1}{4}}\times 2y^{\frac{3}{4}}}{x^2y^{\frac{1}{2}}}\] is what i am reading

OpenStudy (misty1212):

that close?

OpenStudy (anonymous):

Yes exactly

OpenStudy (misty1212):

ok up top we got \(2\times 2=4\) that stays up

OpenStudy (misty1212):

the \(x^{-\frac{1}{4}}\) comes downstairs, as \(x^{\frac{1}{4}}\) and the \(y^{\frac{1}{2}}\) comes upstairs as \(y^{-\frac{1}{2}}\)

OpenStudy (misty1212):

you get \[\huge \frac{4y^{\frac{3}{4}-\frac{1}{2}}}{x^{2+\frac{1}{4}}}\]

OpenStudy (misty1212):

deal with the exponents and you are done it is mostly arithmetic

OpenStudy (anonymous):

oh ok thanks! I have a test tomorrow and I needed to understand it before then

OpenStudy (misty1212):

good luck with the exam